### v1

@@ 2.1 Cluster-Robust Inference
In this paper, we assume that the error terms are independent. Accordingly, we can estimate standard errors that are robust to heteroskedasticity but not robust to clustering. Cluster-robust inference is challenging when models like those in equations (6) and (7) are estimated at the silo level. Researchers typically will cluster at the level of the policy change, which is the silo in this case. Conventional methods of cluster-robust inference will not work with silo-specific data because there is only one cluster in each dataset. Work in progress by the authors of this paper aims to extended these methods to allow for cluster-robust inference.
3 UN–DID When Strong Parallel Trends Hold
In this section, we assume that the strong parallel trends assumption is plausible between the treated and untreated silos. We also assume that no anticipation, homogeneous treatment effects, and no staggered adoption hold. Under these assumptions, we show that the UN–DID and the conventional DID recover the same estimate of the ATT when strong parallel trends hold, both analytically and through a Monte Carlo simulation study.
3.1 Equivalence of the Estimated ATT Between Conventional and UN–DID Methods with No Covariates
@@
Now we will prove that $\hat{\beta_{3}}$ is the sample analogue of Equation (3), or the estimate of the ATT. According to the Frisch–Waugh–Lowell (FWL) theorem, $\hat{\beta_{3}}$ will be the same as the coefficient of $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ from the regression shown in Equation (12). Here, $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ are the residuals from the regression shown in Equation (11).
@@
$\begin{split}\hat{\beta}_{3}=&\sum_{i}\sum_{t}\frac{Y_{i,t}\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)}{\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)^{2}}.\end{split}$ (13)
@@
$\begin{split}&\sum_{i\in t=0,T}Y_{i,0}^{T}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{2}\right)+\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)+\\ &\sum_{i\in t=0,C}Y_{i,1}^{C}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)+\sum_{i\in t=1,C}Y^{C}_{i,1}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{1}\right).\end{split}$
@@
$\begin{split}A=\sum_{i\in t=0,T}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{2}\right)^{2}+\sum_{i\in t=1,T}\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)^{2}+\sum_{i\in t=0,C}\left(-\hat{\alpha}_{0}\right)^{2}+\sum_{i\in t=1,C}\left(-\hat{\alpha}_{0}-\hat{\alpha}_{1}\right)^{2}.\end{split}$
The denominator is the sum of squared residuals from the regression in Equation (11). Combining the numerator and the denominator, $\hat{\beta_{3}}$ can be written as
$\begin{split}\hat{\beta}_{3}=&\frac{\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)\sum_{i\in t=1,T}Y_{i,1}^{T}}{A}-\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)\sum_{i\in t=0,T}Y_{i,0}^{T}}{A}\\ &-\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{1}\right)\sum_{i\in t=1,C}Y_{i,1}^{C}}{A}+\frac{\left(-\hat{\alpha_{0}}\right)\sum_{i\in t=0,C}Y_{i,0}^{C}}{A}.\end{split}$ (14)
In Appendix A, we show that $\frac{\left(1-\hat{\alpha}_{0}-\hat{\alpha}_{1}-\hat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the the treated group and the post-intervention period ($N_{1,1})$ and $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the treated group in the pre-intervention period ($N_{1,0}$). Similarly, $\frac{\left(\hat{\alpha}_{0}+\hat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the control group in the post-intervention period $(N_{0,1})$ and $\frac{\left(-\hat{\alpha}_{0}\right)}{A}$ is reciprocal of the number of observations in the control group in the pre-intervention period ($N_{0,0})$. Therefore, Equation (14) can be further simplified as
@@
UN–DID Regression
In this subsection, we hold that Assumption 6 is violated (data are no longer poolable), while assumptions 1 to 5 hold. Under the violation of Assumption 6, the conventional regression can no longer be used to estimate the ATT. Under these assumptions, we prove that $(\widehat{\lambda_{2}^{T}}-\widehat{\lambda_{1}^{T}})-(\widehat{\lambda_{2}^{C}}-\widehat{\lambda_{1}^{C}})$ derived from the UN–DID regressions can recover the estimate of the ATT. Additionally, we demonstrate that the ATT obtained through the UN–DID regression is numerically equal to the conventional ATT estimate given that the samples used for both approaches are the same. Equations (6) and (7) can be rewritten in the following way:
@@
Comparing equations (18) and (19), we can see that $\lambda_{1}^{j}=\gamma_{0}^{j}$ and $(\lambda_{2}^{j}-\lambda_{1}^{j})=\lambda_{2}^{j}$. Now we will prove that $(\widehat{\lambda_{2}^{T}}-\widehat{\lambda_{1}^{T}})-(\widehat{\lambda_{2}^{C}}-\widehat{\lambda_{1}^{C}})=\hat{\gamma_{1}^{T}}-\hat{\gamma_{1}^{C}}$ is the sample analogue of the ATT shown in Equation (3). In Equation (18), the coefficient of interest is $\hat{\gamma^{j}_{1}}$. According to the FWL theorem, $\hat{\gamma^{j}_{1}}$ will be the same as the coefficient of $(post_{t}^{j}-\widehat{post_{t}^{j}})$ from the regression shown in Equation (21). Here, $(post_{t}^{j}-\widehat{post_{t}^{j}})$ are the residuals from the regression shown in Equation (20).
@@
Similar to in the preceding subsection, we calculate the residuals $(post_{t}^{j}-\widehat{post_{t}^{j}})$ for each observation using Equation (20). For an observation in group j during the pre-intervention period, the residual is given by $(post^{j}_{t}-\widehat{post^{j}_{t}})=-\hat{\eta_{0}^{j}}$. Likewise, for an observation in group j during the post-intervention period, the residual becomes $(post^{j}_{t}-\widehat{post^{j}_{t}})=1-\hat{\eta_{0}^{j}}$. Following this, we proceed to obtain $\widehat{\gamma_{1}^{j}}$ from Equation (21) using the OLS formula:
@@
$\begin{split}\hat{\gamma}_{1}^{i}=&\frac{\sum_{i}\sum_{t}Y_{i,t}^{j}\left(post^{j}_{t}-\hat{post}^{j}_{t}\right)}{\sum_{i}\sum_{t}\left(post^{j}_{t}-\hat{post}^{j}_{t}\right)^{2}}\\ \Rightarrow\hat{\gamma}_{1}^{j}=&\frac{\sum_{i\in t=0}Y_{i,t}^{j}\left(-\hat{\eta}_{0}^{j}\right)+\sum_{i\in t=1}Y_{i,t}^{j}\left(1-\hat{\eta}_{0}^{j}\right)}{\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}\right)^{2}}\\ \Rightarrow\hat{\gamma}_{1}^{j}=&\frac{1-\hat{\eta}_{0}^{j}}{\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}\right)^{2}}\sum_{i\in t=1}Y_{i,1}\\ &-\frac{\hat{\eta}_{0}^{j}}{\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}\right)^{2}}\sum_{i\in t=0}Y_{i,0},\end{split}$
$\hat{\gamma}_{1}^{j}=\frac{1-\hat{\eta}_{0}^{j}}{B}\sum_{i\in t=1}Y_{i,1}-\frac{\hat{\eta}_{0}^{j}}{B}\sum_{i\in t=0}Y_{i,0}.\quad$ (22)
@@
Here, $B=\sum_{i\in t=0}\left(-\hat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\hat{\eta}_{0}^{j}\right)^{2}$. Following the proof in Appendix A, $\frac{1-\hat{\eta}_{0}^{j}}{B}$ is the reciprocal of the number of observations in group $j$ in the post-intervention period and $\frac{\hat{\eta}_{0}^{j}}{B}$ is the reciprocal of the number of observations in group j for the pre-intervention period. Therefore, Equation (22) can be rewritten as
$\hat{\gamma}_{1}^{j}=\frac{1}{N_{j,1}}\sum_{i\in t=1}Y_{i,1}-\frac{1}{N_{j,0}}\sum_{i\in t=0}Y_{i,0}$
$\Rightarrow\hat{\gamma}_{1}^{j}=\left(\overline{Y^{j}_{i,1}}-\overline{Y^{j}_{i,0}}\right).$ (23)
@@
From Equation (23), we can take the difference between $\hat{\gamma^{T}_{1}}$ and $\hat{\gamma^{C}_{1}}$ from the treated and untreated regressions, respectively. Equation (24) shows that the difference is the sample analogue of the ATT shown in Equation (3). If we compare equations (16) and (25), we can see that the conventional estimates and the UN–DID estimates provide numerically equal estimates of the ATT. Note that the conventional regression is not possible with data that are siloed. This is a hypothetical scenario where we assume that we can combine data from different silos together and run both the conventional and the UN–DID regressions on the data. In this case, both methods can identify the ATT and give equivalent estimates:
$\begin{split}\hat{\gamma}_{1}^{T}-\hat{\gamma}_{1}^{C}=\left(\overline{Y^{T}_{i,1}}-\overline{Y^{T}_{i,0}}\right)-\left(\overline{Y^{C}_{i,1}}-\overline{Y^{C}_{i,0}}\right)\end{split}$
@@
3.2 Monte Carlo Design
We conduct a set of Monte Carlo experiments with synthetic data to demonstrate the unbiasedness of the UN–DID estimator and its relative performance compared with the conventional estimator. The overall experiment is depicted in Figure 1, where we apply both conventional and UN–DID regression methods to the same simulated dataset to estimate ATT and assess their properties. Note that, with poolable data, we can apply both the conventional and UN–DID methods to estimate the ATT. With unpoolable data, we are not able to run the conventional regression. Because these are synthetic data, they are poolable. Therefore, we can run both methods on the datasets simultaneously.
Our generated dataset is intended to replicate samples from census data for Ontario and Quebec, specifically concentrating on individuals aged 65 and above. To be more precise, we aligned the age and gender distributions to closely resemble those observed in the populations of Quebec and Ontario in 2001 for individuals aged 65 years and older. We keep the gender constant for all individuals and age them by one year for the next eight years, resulting in our final dataset. In our initial simulation, we created groups of 50 individuals each in both Ontario and Quebec, covering the period from 2000 to 2009. This results in a total of 500 observations evenly distributed over 10 years. We chose a small sample size to ensure the robustness of the results even when dealing with small datasets. For the simulation, we assume that Quebec is treated by some arbitrary policy in the year 2005. This approach ensures a balanced and symmetrical dataset, maintaining an equal number of observations in both pre- and post-treatment periods and in the treatment and control groups. We then repeat the simulations with roughly two thirds of the total observations in Ontario and one third of the total observations in Quebec. So, in a DGP with 500 observations, roughly 333 of the observations are in Ontario and 167 are in Quebec.
Figure 1: Schematic for Simulations
We also consider two data-generating processes to assess whether the estimates of the ATT align between conventional and UN–DID regressions. In the first case, we set the true treatment effect for the treated group at 0. In the second case, the true treatment effect for the treated group is set at 0.1. After generating the data, we perform both conventional and UN–DID regressions to estimate the ATT and access their unbiasedness properties. This is then repeated 1,000 times for each case. The data-generating process for the two cases are as follows:
  • Case 1: No treatment effect, no covariates:
    $Y^{g}_{it}=e^{g}_{it}.$
  • Case 2: Treatment effect, no covariates:
    $Y^{g}_{it}=0.1*DID^{g}_{it}+e^{g}_{it}.$
3.3 Results: No Covariates
In subsection 3.1, we have shown that the estimated ATT between the conventional and the UN–DID regressions are exactly numerically equivalent. To gain a more intuitive understanding of why the estimates from conventional and UN–DID regressions are equal, it is helpful to delve into the interpretation of the coefficients in the UN–DID regressions. Specifically, $\widehat{\lambda_{2}^{T}}$ represents the mean outcome of the treated group in the post-intervention period, while $\widehat{\lambda_{1}^{T}}$ signifies the mean outcome of the treated group in the pre-intervention period. Similarly, $\widehat{\lambda_{2}^{C}}$ corresponds to the mean outcome of the control group in the post-intervention period and $\widehat{\lambda_{1}^{C}}$ denotes the mean outcome of the control group in the pre-intervention period.
@@ 3.3 Results: No Covariates
Therefore, the estimate of $\lambda_{1}^{T}$ is the mean outcome of the treated group in the pre-intervention period. Taking the expectation conditional on $post_{t}^{T}=1$ on both sides, we can also show that $E[Y^{T}_{i,t}|post_{t}^{T}=1]=\lambda_{2}^{T}$. The same can be shown for $\lambda_{1}^{C}$ and $\lambda_{2}^{T}$.
A more formal proof of the interpretations is shown in subsection 3.1. Therefore, $(\widehat{\lambda_{2}^{T}}-\widehat{\lambda_{1}^{T}})-(\widehat{\lambda_{2}^{C}}-\widehat{\lambda_{1}^{C}})$ is an estimate of the ATT shown in Equation (3). In subsection 3.1, we also establish that $\widehat{\beta_{3}}=\left(\overline{Y}_{D_{i}=1,P_{t}=1}-\overline{Y}_{D_{i}=1,P_{t}=0}\right)-\left(\overline{Y}_{D_{i}=0,P_{t}=1}-\overline{Y}_{D_{i}=0,P_{t}=0}\right)$. Here, $\overline{Y}_{D_{i}=k,P_{t}=l}$ represents the mean outcome of group $k$ in period $l$. When the law of large numbers holds, $\widehat{\beta_{3}}$ serves as the estimate of the ATT, as indicated in Equation (3). Assuming the sample is identical for both the conventional and UN–DID methods, both estimators should yield equal estimates of the ATT and their associated standard errors.

### v2

@@ 2.1 Cluster-Robust Inference
In this paper, we assume that the error terms are independent. Accordingly, we can estimate standard errors that are robust to heteroskedasticity but not robust to clustering. It is straightforward to extend the procedures that we discuss to clustering at the sub-silo level. For instance, one could cluster at the person-level when using panel data, or at the city level if the silos are state level. However, researchers will typically cluster at the level of the policy change (Bertrand et al., 2004; MacKinnon et al., 2023), which we have assumed is by silo. Silo level clustering is challenging when models like those in equations (6) and (7) are also estimated at the silo level. Conventional methods of cluster-robust inference will not work with silo-specific data because there is only one cluster in each dataset. With only one cluster the variance matrix is rank deficient and cannot be computed. Work in progress by the authors of this paper aims to extended these methods to allow for silo-level clustering using alternative inference procedures.
3 UN-DID When Strong Parallel Trends Hold
In this section, we assume that the strong parallel trends assumption is plausible between the treated and untreated silos. We also assume that no anticipation, homogeneous treatment effects, and no staggered adoption hold. Under these assumptions, we show that the UN-DID and the conventional DID recover the same estimate of the ATT, both analytically and through a Monte Carlo simulation study.
3.1 Equivalence of the Estimated ATT Between Conventional and UN-DID Methods with No Covariates
@@
Now we will show that $\hat{\beta_{3}}$ is the sample analogue of Equation (3), or the estimate of the ATT. According to the Frisch–Waugh–Lowell (FWL) theorem, $\hat{\beta_{3}}$ will be the same as the coefficient of $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ from the regression shown in Equation (12). Here, $(P_{t}*D_{i}-\widehat{P_{t}*D_{i}})$ are the residuals from the regression shown in Equation (11).
@@
$\begin{split}\widehat{\beta}_{3}=&\sum_{i}\sum_{t}\frac{Y_{i,t}\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)}{\left(P_{t}*D_{i}-\widehat{P_{t}*D_{i}}\right)^{2}}.\end{split}$ (13)
@@
$\begin{split}&\sum_{i\in t=0,T}Y_{i,0}^{T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}\right)+\sum_{i\in t=1,T}Y_{i,1}^{T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)+\\ &\sum_{i\in t=0,C}Y_{i,1}^{C}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)+\sum_{i\in t=1,C}Y^{C}_{i,1}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}\right).\end{split}$
@@
$\begin{split}A=\sum_{i\in t=0,T}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{2}\right)^{2}+\sum_{i\in t=1,T}\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)^{2}+\sum_{i\in t=0,C}\left(-\widehat{\alpha}_{0}\right)^{2}+\sum_{i\in t=1,C}\left(-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}\right)^{2}.\end{split}$
The denominator is the sum of squared residuals from the regression in Equation (11). Combining the numerator and the denominator, $\widehat{\beta}_{3}$ can be written as
$\begin{split}\widehat{\beta}_{3}=&\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)\sum_{i\in t=1,T}Y_{i,1}^{T}}{A}-\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)\sum_{i\in t=0,T}Y_{i,0}^{T}}{A}\\ &-\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{1}\right)\sum_{i\in t=1,C}Y_{i,1}^{C}}{A}+\frac{\left(-\widehat{\alpha}_{0}\right)\sum_{i\in t=0,C}Y_{i,0}^{C}}{A}.\end{split}$ (14)
In Appendix A, we show that $\frac{\left(1-\widehat{\alpha}_{0}-\widehat{\alpha}_{1}-\widehat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the treated group and the post-intervention period ($N_{1,1})$ and $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the treated group in the pre-intervention period ($N_{1,0}$). Similarly, $\frac{\left(\widehat{\alpha}_{0}+\widehat{\alpha}_{2}\right)}{A}$ is the reciprocal of the number of observations in the control group in the post-intervention period $(N_{0,1})$ and $\frac{\left(-\widehat{\alpha}_{0}\right)}{A}$ is reciprocal of the number of observations in the control group in the pre-intervention period ($N_{0,0})$. Therefore, Equation (14) can be further simplified as
@@
UN-DID Regression
In this subsection, we hold that Assumption 6 is violated (data are no longer poolable), while assumptions 1 to 5 hold. Under the violation of Assumption 6, the conventional regression can no longer be used to estimate the ATT. Under these assumptions, we prove that $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})$ derived from the UN-DID regressions can recover the estimate of the ATT. Provided that the data samples used for both approaches are the same, we demonstrate that the ATT obtained through the UN-DID regression is numerically equal to the conventional ATT estimate. Equations (6) and (7) can be rewritten in the following way:
@@
Comparing equations (18) and (19), we can see that $\lambda_{1}^{j}=\gamma_{0}^{j}$ and $(\lambda_{2}^{j}-\lambda_{1}^{j})=\lambda_{2}^{j}$. Now we will prove that $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})=\hat{\gamma}_{1}^{T}-\hat{\gamma}_{1}^{C}$ is the sample analogue of the ATT shown in Equation (3). In Equation (18), the coefficient of interest is $\hat{\gamma}^{j}_{1}$. According to the FWL theorem, $\hat{\gamma^{j}_{1}}$ will be the same as the coefficient of $(post_{t}^{j}-\widehat{post}_{t}^{j})$ from the regression shown in Equation (21). Here, $(post_{t}^{j}-\widehat{post}t^{j})$ are the residuals from the regression shown in Equation (20).
@@
Similar to in the preceding subsection, we calculate the residuals $(post_{t}^{j}-\widehat{post_{t}^{j}})$ for each observation using Equation (20). For an observation in group j during the pre-intervention period, the residual is given by $(post^{j}_{t}-\widehat{post^{j}_{t}})=-\widehat{\eta_{0}^{j}}$. Likewise, for an observation in group j during the post-intervention period, the residual becomes $(post^{j}_{t}-\widehat{post^{j}_{t}})=1-\widehat{\eta_{0}^{j}}$. Following this, we proceed to obtain $\widehat{\gamma}_{1}^{j}$ from Equation (21) using the OLS formula:
@@
$\begin{split}\widehat{\gamma}_{1}^{i}=&\frac{\sum_{i}\sum_{t}Y_{i,t}^{j}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)}{\sum_{i}\sum_{t}\left(post^{j}_{t}-\widehat{post}^{j}_{t}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{\sum_{i\in t=0}Y_{i,t}^{j}\left(-\widehat{\eta}_{0}^{j}\right)+\sum_{i\in t=1}Y_{i,t}^{j}\left(1-\widehat{\eta}_{0}^{j}\right)}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}}\\ \Rightarrow\widehat{\gamma}_{1}^{j}=&\frac{1-\widehat{\eta}_{0}^{j}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}}\sum_{i\in t=1}Y_{i,1}\\ &-\frac{\widehat{\eta}_{0}^{j}}{\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}}\sum_{i\in t=0}Y_{i,0},\end{split}$
$\widehat{\gamma}_{1}^{j}=\frac{1-\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=1}Y_{i,1}-\frac{\widehat{\eta}_{0}^{j}}{B}\sum_{i\in t=0}Y_{i,0}.\quad$ (22)
@@
Here, $B=\sum_{i\in t=0}\left(-\widehat{\eta}_{0}^{j}\right)^{2}+\sum_{i\in t=1}\left(1-\widehat{\eta}_{0}^{j}\right)^{2}$. Following the proof in Appendix A, $\frac{1-\widehat{\eta}_{0}^{j}}{B}$ is the reciprocal of the number of observations in group $j$ in the post-intervention period and $\frac{\widehat{\eta}_{0}^{j}}{B}$ is the reciprocal of the number of observations in group j for the pre-intervention period. Therefore, Equation (22) can be rewritten as:
$\widehat{\gamma}_{1}^{j}=\frac{1}{N_{j,1}}\sum_{i\in t=1}Y_{i,1}-\frac{1}{N_{j,0}}\sum_{i\in t=0}Y_{i,0}$
$\Rightarrow\widehat{\gamma}_{1}^{j}=\left(\overline{Y}^{j}_{i,1}-\overline{Y}^{j}_{i,0}\right).$ (23)
@@
From Equation (23), we can take $\widehat{\gamma}^{T}_{1}$ and $\widehat{\gamma}^{C}_{1}$ from the treated and untreated regressions, respectively, and then take the difference between them. Equation (24) shows that the difference is the sample analogue of the ATT shown in Equation (3). If we compare equations (16) and (25), we can see that the conventional DID estimates and the UN-DID estimates provide numerically equal estimates of the ATT. Note that the conventional regression is not possible with data that are siloed. In a hypothetical scenario, we can combine data from different silos together and run both the conventional and the UN-DID regressions on the data. In this case, both methods can identify the ATT and give equivalent estimates:
$\begin{split}\widehat{\gamma}_{1}^{T}-\widehat{\gamma}_{1}^{C}=\left(\overline{Y}^{T}_{i,1}-\overline{Y}^{T}_{i,0}\right)-\left(\overline{Y}^{C}_{i,1}-\overline{Y}^{C}_{i,0}\right)\end{split}$
@@
To gain a more intuitive understanding of why they are equal, it is helpful to delve into the interpretation of the coefficients in the UN-DID regressions. Specifically, $\widehat{\lambda_{2}^{T}}$ represents the mean outcome of the treated group in the post-intervention period, while $\widehat{\lambda_{1}^{T}}$ signifies the mean outcome of the treated group in the pre-intervention period. Similarly, $\widehat{\lambda_{2}^{C}}$ corresponds to the mean outcome of the control group in the post-intervention period and $\widehat{\lambda_{1}^{C}}$ denotes the mean outcome of the control group in the pre-intervention period.
@@
Therefore, the estimate of $\lambda_{1}^{T}$ is the mean outcome of the treated group in the pre-intervention period. Taking the expectation conditional on $post_{t}^{T}=1$ on both sides, we can also show that $E[Y^{T}_{i,t}|post_{t}^{T}=1]=\lambda_{2}^{T}$. The same can be shown for $\lambda_{1}^{C}$ and $\lambda_{2}^{T}$. Based on this result, we can say that $(\widehat{\lambda}_{2}^{T}-\widehat{\lambda}_{1}^{T})-(\widehat{\lambda}_{2}^{C}-\widehat{\lambda}_{1}^{C})$ is an estimate of the ATT shown in Equation (3). We also establish that $\widehat{\beta}_{3}=\left(\overline{Y}_{D_{i}=1,P_{t}=1}-\overline{Y}_{D_{i}=1,P_{t}=0}\right)-\left(\overline{Y}_{D_{i}=0,P_{t}=1}-\overline{Y}_{D_{i}=0,P_{t}=0}\right)$ from Equation (10). Here, $\overline{Y}_{D_{i}=k,P_{t}=l}$ represents the mean outcome of group $k$ in period $l$. When the law of large numbers holds, $\widehat{\beta}_{3}$ serves as the estimate of the ATT, as indicated in Equation (3). Assuming the sample is identical for both the conventional and UN-DID methods, both estimators should yield equal estimates of the ATT and their associated standard errors.
3.2 Monte Carlo Design
We conduct a set of Monte Carlo experiments with synthetic data to demonstrate the unbiasedness of the UN-DID estimator and its relative performance compared with the conventional DID estimator. The overall experiment is depicted in Figure 1, where we apply both conventional and UN-DID regression methods to the same simulated dataset to estimate ATTs and assess their properties. Note that, with poolable data, we can apply both the conventional and UN-DID methods to estimate the ATT. With unpoolable data, we are not able to run the conventional regression. Because these are synthetic data, they are poolable. Therefore, we can run both methods on the dataset.
Our generated dataset is intended to replicate samples from census data for Ontario and Quebec, specifically concentrating on individuals aged 65 and above. To be more precise, we aligned the age and gender distributions to closely resemble those observed in the populations of Quebec and Ontario in 2001 for individuals aged 65 years and older. We keep the gender constant for all individuals and age them by one year for the next nine years, resulting in our final dataset. Our outcome is a continuous variable. In our initial simulation, we created groups of 50 individuals each in both Ontario and Quebec, covering the period from 2000 to 2009. This results in a total of 500 observations evenly distributed over 10 years. We chose a small sample size to ensure the robustness of the results even when dealing with small datasets. For the simulation, we assume that Quebec is treated by some arbitrary policy in the year 2005. This approach ensures a balanced and symmetrical dataset, maintaining an equal number of observations in both pre- and post-treatment periods and in the treatment and control groups. We then repeat the simulations with roughly two thirds of the total observations in Ontario and one third of the total observations in Quebec. So, in a DGP with 500 observations, roughly 333 of the observations are in Ontario and 167 are in Quebec.
Figure 1: Schematic for Simulations
We also consider two data-generating processes to assess whether the estimates of the ATT align between conventional and UN-DID regressions. In the first case, we set the true treatment effect for the treated group at 0. In the second case, the true treatment effect for the treated group is set at 0.1. After generating the data, we perform both conventional and UN-DID regressions to estimate the ATT and access their unbiasedness properties. This is then repeated 1,000 times for each case. The data-generating process for the two cases are as follows:
  • Case 1: No treatment effect, no covariates:
    $Y^{g}_{it}=e^{g}_{it}.$
  • Case 2: Treatment effect, no covariates:
    $Y^{g}_{it}=0.1*DID^{g}_{it}+e^{g}_{it}.$
In this simulation study, the generated dataset is a balanced panel. The results displayed in the ensuring sections also apply to repeated cross sectional data. The results for time invariant covariate with repeated cross sectional data are comparable to the results with time varying covariates with panel data, as the distribution of the covariate varies over time due to different individualsbeing present in the data in each year.
3.3 Results: No Covariates